# Embedding Space Allocation with Angle-Norm Joint Classifiers for Few-Shot Class-Incremental Learning

### v1

@@ 5.4 Numerical Analysis
Visualization of Embedding Separation. We randomly choose 6 classes from base classes and 3 classes from novel classes respectively and visualize the embeddings of them on CIFAR100 dataset with t-SNE [57] in Fig. 5. Firstly, it can be observed that the relative positions of class embeddings shift compared to Decoupled-Cosine. In Decoupled-Cosine, the purple, brown, and red classes are adjacent, while in SAAN, the green, black, and red classes are adjacent. This is because the allocation of class centers affects the positions of class means. During the class increment process, in SAAN, the purple and brown new classes are assigned to two different reserved spaces, whereas in Decoupled-Cosine, the purple and brown classes are assigned to adjacent spaces and are very close to each other. This makes it difficult to distinguish between the purple and brown classes in Decoupled-Cosine. While in SAAN, it is easier to identify them. Additionally, in SAAN, the inter-class distance is larger, while the intra-class distance is smaller.
In Fig. 5(b), it can be observed that the sample embeddings are distributed around the centers we set, indicating that the limiting effect of the centers is effective. There is some deviation between these centers and the class mean, which is due to the influence of CE.
@@
Performance on Novel and Base Classes. Solely based on the accuracy over all classes, we cannot evaluate how much new knowledge the model has learned and how much old knowledge it has forgotten. We report the accuracy of the base classes and the novel classes as well as harmonic mean (HM) of them in Fig. 6. With the help of SAAN, SAVC’s accuracy for novel classes improves by 7.20% and 3.76% in CIFAR100 and CUB200 respectively. The accuracy improvement for old classes is very limited compared to that for new classes. SAAN obviously enhances the compatibility of the model to new classes verifying that limiting the space occupation of old class and allocating space for new class are conducive to the learning of new classes.
Figure 7: Last session accuracy of SAVC + SAAN with different hyper-parameters. (a) $\alpha$ varies while $\eta$ is fixed. (b) $\eta$ varies while $\alpha$ is fixed.
Parameter Sensitivity. We maintain a fixed ratio of 5 between $\alpha$ and $\beta$ and vary $\alpha$ and $\eta$ to access the impact of hyper-parameters. Fig. 7(a) demonstrates that $\alpha$ is highly sensitive. As $\alpha$ increases, the model performance first increases and then decreases. This is because if $\alpha$ is too small, the effect of SAAN will be limited. If $\alpha$ is too large, the sample points will overfit the centers. In contrast, $\eta$ is not as sensitive. Fortunately, any $\eta$ ranging from 0.05 to 1 can significantly improve the performance. However, a very small value of $\eta$ approximates a pre-set and fixed prototype, while a very large $\eta$ results in unstable centers. Both cases are unfavorable for spatial allocation and guiding sample learning.
@@
Angle Distribution from the Centers. We present the angle of each sample relative to its center of SAVC and SAVC+SAAN, as shown in Fig. 8. It can be observed that the mean angle for novel classes of SAVC is approximately 60 degrees, while the mean angle of SAVC+SAAN is around 50 degrees, with a noticeably smaller maximum angle compared to SAVC. The results on base classes are similar. This implies that SAVC+SAAN exhibits a more compact intra-class structure, by using our CC loss.
Figure 9: Quantile-Quantile plot of embedding log-norms on miniImageNet dataset.
Normal Test. We present the Quantile-Quantile plot of the log-norms of smaple embeddings in Fig. 9, where the horizontal axis represents theoretical quantities and the vertical axis represents sample quantities. Due to the granularity of the base session distribution being categories, we randomly select one category. During the incremental sessions, we select one session due to distribution sharing. It can be observed that the data points are nearly aligned along a straight line, indicating that the sample log-norms can be considered to follow a normal distribution. In session 8, the data points slightly deviate above the straight line, indicating that the shared norm distribution exhibits a long-tail behavior.
@@ 6 Conclusion
Future Works. In the current FSCIL problem setting, the number of future classes is known in advance, which facilitates prior spatial division. However, in more open incremental scenarios where future classes are uncertain, strategies for initial spatial allocation and distribution require further exploration. The log-normal distribution assumption for the norm may not be sufficiently accurate, especially in the shared distribution during the incremental sessions. Therefore, there may be better assumptions or estimation methods for the distribution.
@@ 7 Acknowledgments
Appendix A Gradient Analysis of CC Loss
In this section, we provide the gradient of the proposed CC loss and analyze the update mechanism of the learnable centers. Since analyzing the gradients of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ with respect to embeddings requires deriving the gradient of the cosine similarity with respect to the vector, we give it in advance:
@@ Appendix A Gradient Analysis of CC Loss
where $\mathbf{a}$ and $\mathbf{b}$ are any venctors. In this paper, the norm specifically refers to the L2-norm.
@@ Appendix B Comparison Methods
iCaRL [39]. It maintains an ”episodic memory” of the exemplars, which enables to incrementally learn new classes without forgetting old classes.
@@
SAVC [16]. It introduces more semantic knowledge by imagining virtual classes which not only hold enough feature space for future updates, but enable a multi semantic aggregated inference effect.
Appendix C Repeated Experiments on 2SNCM
Table 5: Comparison 2SNCM with NCM on three mainstream datasets in FSCIL. $\Delta_{avg}$: The average of the accuracies on all session. The data comes from the average accuracy of 5 random experiments.
@@ Appendix C Repeated Experiments on 2SNCM
2SNCM and NCM only differ in the selection of representative points. In order to exclude random perturbations, we conducted multiple repeated experiments on them. Here the seeds are canceled and multiple repeated experiments are conducted. We use Decoupled-Cosine to conduct 5 replicate experiments on 3 datasets respectively, and report the average performance of them in Tab. 5. Compared with NCM, 2SNCM improve the average accuraciers on CIFAR100, CUB200 by 0.16%. On miniImageNet, 2SNCM performs better in the early incremental phase and base phase, and NCM performs better in the late incremental phase. The better performance of 2SNCM indicates the effectiveness of normalizing based on angles when there are enough base round samples, while preserving norm information without normalization when the incremental round samples are fewer.

### v2

@@ 5.5 Numerical Analysis
Visualization of Embedding Separation. We randomly choose 6 classes from base classes and 3 classes from novel classes respectively and visualize the embeddings of them on the CIFAR100 dataset with t-SNE [62] in Fig. 5. Firstly, it can be observed that the relative positions of class embeddings shift compared to Decoupled-Cosine. In Decoupled-Cosine, the purple, brown, and red classes are adjacent, while in SAAN, the green, black, and red classes are adjacent. This is because the allocation of class centers affects the positions of class means. During the class increment process, in SAAN, the purple and brown new classes are assigned to two different reserved spaces, whereas in Decoupled-Cosine, the purple and brown classes are assigned to adjacent spaces and are very close to each other. This makes it difficult to distinguish between the purple and brown classes in Decoupled-Cosine. While in SAAN, it is easier to identify them. Additionally, in SAAN, the inter-class distance is larger, while the intra-class distance is smaller.
In Fig. 5(b), it can be observed that the sample embeddings are distributed around the centers we set, indicating that the limiting effect of the centers is effective. There is some deviation between these centers and the class mean, which is due to the influence of cross-entropy.
@@
Performance on Novel and Base Classes. Solely based on the accuracy over all classes, we cannot evaluate how much new knowledge the model has learned and how much old knowledge it has forgotten. We report the accuracy of the base classes and the novel classes as well as harmonic mean in Fig. 6. With the help of SAAN, SAVC’s accuracy for novel classes improves by 7.20% and 3.76% in CIFAR100 and CUB200 respectively. The accuracy improvement for old classes is very limited compared to that for new classes. SAAN obviously enhances the compatibility of the model to new classes verifying that limiting the space occupation of old classes and allocating space for new classes are conducive to the learning of new classes.
Figure 7: The parameter grid search for the last session accuracy of SAVC + SAAN on the CUB200 dataset. (a) The weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. (b) The center moving rate and its decay rate.
Figure 8: Last session accuracy of SAVC + SAAN with different hyper-parameters on three datasets. (a) $\alpha$ varies while $\eta$ is fixed and the ratio of $\alpha$ to $\beta$ is 5. (b) $\eta$ varies while $\alpha$ is fixed, and $\lambda$ is 0.1. (c) $C$ varies.
Parameter Sensitivity. We conduct a search and discussion on the loss weights $\alpha$, $\beta$, compression coefficient $C$, as well as the center moving rate $\eta$ and its decay rate $\lambda$. The grid search results for cosine center loss weights and center moving rate with its decay rate on CUB200 are shown in Fig. 7. The parameter analysis across three datasets is presented in Fig. 8. As $\alpha$ and $\beta$ increase, the effect of CCSA becomes stronger, leading to improved model performance, which reaches its optimum when $\alpha$ is 2 and $\beta$ is 0.4. $\mathcal{L}_{1}$ serves to push away from other centers, while $\mathcal{L}_{2}$ pulls closer to the corresponding center. If $\beta$ is too large and $\alpha$ is too small, the pushing effect will be dominant, while the more crucial pulling effect will be weakened, leading to a negative impact. For all three datasets, the optimal value of $C$ is consistently 0.005. When $C$ exceeds 0.05, the influence of norm logits becomes overly dominant, rendering the more important angular logits ineffective, which leads to a sharp performance drop.
It can be observed that the center moving rate and decay rate are relatively insensitive, as values ranging from 0.25 to 2 all have a positive effect. However, considering two special cases: when the center moving rate is 0, it represents fixed centers; when the center moving decay rate is 1, it means that the centers are continuously updated without playing a role in spatial allocation, only contributing to intra-class cohesion. Fixed centers perform better than the baseline but are inferior to the learnable SAAN. This indicates that spatial allocation is beneficial for FSCIL, and learnable centers achieve better allocation performance. Continuously updating centers perform almost the same as the baseline, failing to achieve spatial allocation. This highlights the importance of spatial allocation, as merely increasing inter-class distance has limited effectiveness.
@@
Angle Distribution from the Centers. We present the angle of each sample relative to its center of SAVC and SAVC+SAAN, as shown in Fig. 9. It can be observed that the mean angle for novel classes of SAVC is approximately 60 degrees, while the mean angle of SAVC+SAAN is around 50 degrees, with a noticeably smaller maximum angle compared to SAVC. The results on base classes are similar. This implies that SAVC+SAAN exhibits a more compact intra-class structure, by using our cosine center loss.
Figure 10: Quantile-Quantile plot of embedding log norms on miniImageNet dataset.
Normal Test.
We present the Quantile-Quantile plot of the log norms of sample embeddings in Fig. 10, where the horizontal axis represents theoretical quantities and the vertical axis represents sample quantities. Due to the granularity of the base session distribution being categories, we randomly select one category. During the incremental sessions, we select one session due to distribution sharing. It can be observed that the data points are nearly aligned along a straight line, indicating that the sample log norms can be considered to follow a normal distribution. In session 8, the data points slightly deviate above the straight line, indicating that the shared norm distribution exhibits a long-tail behavior.
@@ 7 Acknowledgments
Appendix A Gradient Analysis of Cosine Center Loss
In this section, we provide the gradient of the proposed cosine center loss and analyze the update mechanism of the learnable centers. Since analyzing the gradients of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ with respect to embeddings requires deriving the gradient of the cosine similarity with respect to the vector, we give it in advance:
@@ Appendix A Gradient Analysis of Cosine Center Loss
where $\mathbf{a}$ and $\mathbf{b}$ are any venctors. In this paper, the norm specifically refers to the L2 norm.
@@ Appendix B Comparison Methods
iCaRL [41]. It maintains an ”episodic memory” of the exemplars, which enables one to incrementally learn new classes without forgetting old classes.
@@
SAVC [16]. It introduces more semantic knowledge by imagining virtual classes which not only hold enough feature space for future updates but enable a multi-semantic aggregated inference effect.
Appendix C Detailed Results of the Open-Ended Experiments
To better simulate the real-world environment of class-incremental tasks, we conduct this comparative experiment under open-ended and imbalanced conditions. The experimental setup and the analysis are elaborated in Sec. 5.3. The experimental results we obtain are presented in Tab. 7.
Table 7: Detailed results of the open-ended and imbalanced experiments on CompCars dataset.
Method Random Acc. in each session (%$\uparrow$
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
Decoupled-Cosine[60] rand 1 85.88 83.90 82.88 81.79 80.57 79.13 76.44 73.40 71.61 69.15 67.27 64.40 63.34 61.57 60.58 59.77 58.71 57.87 57.44 56.30 56.22
rand 2 86.26 85.02 83.89 81.68 80.69 79.15 76.53 74.11 71.93 69.49 67.27 64.80 63.56 61.95 60.94 59.02 58.58 57.72 56.96 56.37 55.56
rand 3 85.88 84.36 83.16 82.21 80.85 79.46 77.00 74.30 72.37 70.07 67.90 65.29 64.10 62.32 61.30 60.50 59.52 58.53 58.10 57.38 56.96
SAAN rand 1 93.79 91.89 90.70 89.80 88.51 86.98 84.44 81.53 79.79 77.41 75.16 72.23 71.24 69.12 68.15 67.06 66.05 64.96 64.33 63.37 62.91
rand 2 93.40 91.80 90.45 88.83 87.77 85.96 83.08 80.40 78.36 75.57 73.44 70.78 69.55 67.75 66.56 64.62 64.09 63.15 62.29 61.56 60.82
rand 3 93.79 91.95 90.97 90.07 88.78 87.31 84.83 81.97 80.26 77.58 75.49 72.58 71.35 69.28 68.09 67.05 66.06 65.01 64.39 63.50 63.09
FACT[11] rand 1 92.76 91.00 89.61 88.74 87.18 85.73 83.10 80.07 78.25 75.64 73.23 71.03 69.55 67.72 66.71 65.74 64.75 63.80 63.30 62.48 62.10
rand 2 92.82 91.40 89.59 87.96 86.97 85.24 82.64 79.98 77.93 75.18 72.91 70.31 68.96 67.08 66.00 63.97 63.54 62.67 61.83 61.16 60.15
rand 3 92.76 90.85 89.81 88.87 87.43 85.93 83.31 80.44 78.79 76.17 73.76 71.56 70.06 68.07 67.00 66.06 65.13 64.05 63.54 62.83 62.40
FACT+SAAN rand 1 93.44 91.44 89.77 88.92 87.87 86.27 84.05 81.41 79.64 77.16 75.21 72.90 71.48 69.49 68.49 67.75 66.92 66.00 65.54 64.78 64.40
rand 2 92.54 90.72 89.53 88.05 87.07 85.26 82.76 80.20 78.23 75.48 73.70 71.06 69.83 68.02 66.92 64.96 64.55 63.62 62.76 62.26 61.46
rand 3 93.61 92.19 90.65 89.85 88.70 87.05 84.89 82.48 80.70 78.43 76.02 73.74 72.32 70.32 69.37 68.32 67.50 66.49 65.97 65.56 64.92
SAVC[16] rand 1 95.31 93.18 91.92 91.10 90.53 89.01 87.43 85.07 84.05 81.91 80.29 78.37 77.59 76.22 75.47 74.98 74.20 73.35 73.11 72.09 71.85
rand 2 94.38 92.37 91.45 90.39 89.11 87.78 86.11 84.20 82.53 80.34 78.65 76.71 75.38 73.93 73.11 71.36 71.05 70.32 69.44 68.82 68.10
rand 3 95.31 93.69 92.42 91.75 90.97 89.77 88.16 86.08 85.06 82.98 81.54 79.36 78.33 76.81 76.11 75.36 74.51 73.84 73.36 72.93 72.33
SAVC+SAAN rand 1 95.60 94.13 92.86 92.52 91.74 90.40 88.73 86.63 85.77 83.91 83.02 80.95 79.94 78.59 77.74 77.36 76.63 75.71 75.45 74.56 74.37
rand 2 95.35 94.08 93.52 92.25 91.07 89.83 87.96 86.36 84.87 82.71 81.30 79.81 78.89 77.37 76.63 75.10 74.58 73.78 72.88 72.37 71.86
rand 3 95.60 94.63 93.46 92.79 91.91 90.68 89.18 87.32 86.23 84.37 83.02 81.12 80.18 78.74 77.78 77.20 76.46 75.67 75.27 74.57 74.53
Appendix D Computational Cost Analysis
Table 8: Analysis of trainable parameter count and inference computational cost for different methods on the CUB200 dataset. The computational cost is measured with a batch size of 100.
Method Model Parameters (M) Inference Cost (GFLOPs) Center Parameters (M) ANJ Cost (GFLOPs)
Decoupled-Cosine[60] 11.18 182.35 +0.0512 +1.68
FACT[11] 11.18 182.35 +0.0512 +1.68
SAVC [16] 11.50 364.77 +0.1024 +3.27
We analyze the computational cost of SAAN compared to other methods on CUB200 dataset. The additional cost introduced by SAAN mainly comes from the increased parameters of trainable centers and the inference computation brought by ANJ. Therefore, we report the extra training parameters and inference computation introduced by SAAN, as well as the parameter count and inference computation of other methods themselves, as shown in Tab. 8. It can be observed that the additional parameter count and inference computation introduced by SAAN are both less than 1%. In SAVC, the computational cost and center parameter count are both twice as high as in other methods because SAVC performs category augmentation with virtual classes, leading to a doubled number of centers. During inference, both the original and augmented samples need to be processed, resulting in increased inference computation.
Appendix E Repeated Experiments on 2SNCM
Table 9: Comparison 2SNCM with NCM on three mainstream datasets in FSCIL. $\Delta_{avg}$: The average of the accuracies on all sessions. The data comes from the average accuracy of 5 random experiments.
@@ Appendix E Repeated Experiments on 2SNCM
2SNCM and NCM only differ in the selection of representative points. In order to exclude random perturbations, we conducted multiple repeated experiments on them. Here the seeds are canceled and multiple repeated experiments are conducted. We use Decoupled-Cosine to conduct 5 replicate experiments on 3 datasets respectively and report the average performance of them in Tab. 9. Compared with NCM, 2SNCM improves the average accuracies on CIFAR100, CUB200 by 0.16%. On miniImageNet, 2SNCM performs better in the early incremental phase and base phase, and NCM performs better in the late incremental phase. The better performance of 2SNCM indicates the effectiveness of normalizing based on angles when there are enough base round samples while preserving norm information without normalization when the incremental round samples are fewer.